# The Harrow-Hassidim-Lloyd algorithm with qutrits

### v1

@@ Abstract
We extend the Harrow-Hassidim-Lloyd (HHL) algorithm, which is well-studied in the qubit framework, to its qutrit counterpart (which we call qutrit HHL, as opposed to qubit HHL, which is HHL using qubits). We design the circuit for the algorithm and develop a program for its implementation. We test HHL with qutrits for simple matrices and verify the results against the expected outcomes. We apply the algorithm to quantum chemistry, and in particular, to the potential energy curve calculations of the model problem of the hydrogen molecule in the split valence basis. We compare the number of qudits and the number of gates required between qubit and qutrit HHL implementations. In general, we find that for a fixed precision, the qutrit HHL circuit requires fewer number of qudits and comparable number of two-qudit gates than its qubit counterpart.
@@ I Introduction
All of the above discussed works involve qubits. Recent works have identified that using qudits for quantum algorithms can have potential advantages over their qubit counterparts [qudit-adv, qutrit-benchmarking]. Several algorithms like Deutsch-Josza, Bernstein-Vazirani, quantum Fourier transform (QFT), QPE, etc, have been discussed in the qudit framework [qudit-algorithms]. Despite the large volume of ongoing research in quantum algorithms and computing, there remains a relative paucity of literature on higher-dimensional qudits, and in particular, qutrits, where a qutrit has an additional quantum level relative to a qubit, but which is a natural extension to consider. Furthermore, in view of some of the recent advances in qutrit hardware [qutrit-hardware, qutrit-hardware-gates, qutrit-benchmarking], it is timely to explore the domain of quantum algorithms with qutrits. Therefore, given the potential of HHL for offering advantage across a wide variety of problems and the timely need to investigate quantum algorithms using qutrits, we extend the HHL algorithm to the qutrit framework (which we shall term as qutrit HHL hereafter). Furthermore, we consider quantum chemistry as a pilot application for qutrit HHL and to that end, carry out calculations to generate the potential energy curve for the model system of $H_{2}$ in a split valence basis. This is followed by a comparative analysis of the costs involved between qubit and qutrit HHL calculations. To the best of our knowledge, one previous work proposes qutrit HHL, but only presents a preliminary circuit design [old-qutrit-hhl].
The subsequent sections are organized as follows: Section II introduces the HHL algorithm for qubits. Following a primer on ternary logic and qutrit gates in Sections III.1 and III.2 respectively, we move to our qutrit HHL implementation is discussed in Section III.3. We then provide a brief introduction to the quantum chemistry application, in which we solve for the correlation energy using HHL (Section IV). We then discuss results from our numerical simulations in Section V, which include toy matrices (Section V.1) and quantum chemistry (Section V.2). We then carry out a comparative analysis between qubits and qutrits in Section VI, where we discuss estimates for the number of qudits required (Section VI.1) and the number of two-qudit gates incurred for HHL calculations in the qubit and qutrit cases (Section VI.2). We finally conclude in Section VII.
@@ II The HHL algorithm
The HHL algorithm is discussed in detail, for example, in Reference [HHL_walkthru], but here we provide only the relevant details that we would need for a subsequent qutrit extension of the algorithm. Figure 1 gives a schematic quantum circuit representation of the algorithm. The $N\crossproduct 1$ dimensional vector, $\vec{b}$, which is an input to the algorithm, is encoded in an $m$-dimensional normalized quantum state $|b\rangle_{m}$ in the state register via amplitude encoding such that $2^{m}=N$ for qubits. On the other hand, the matrix $A$ is used in the circuit as a part of a controlled-unitary of the form $CU=C\ e^{iAt}$, where $U=e^{iAt}$ for an appropriately chosen scalar, $t$. As the figure shows, the HHL circuit broadly comprises of three parts - (a) QPE, (b) Controlled-rotation, and (c) Inverse of QPE. These steps are followed by measurement and post-selection. The above-mentioned controlled-unitaries, which are controlled on a set of $n_{r}$ clock register qubits, are used for quantum phase estimation. At the end of the QPE module, the eigenvalues of $A$ are stored in the clock register with a precision of $n_{r}$ digits. The parametrized controlled-rotation module, $CR(\theta)$, implements inversion of eigenvalues using an ancillary qubit that we call the HHL ancilla for brevity. The inverse QPE module is then applied in order to disentangle the clock register from the rest of the system. Post-selection of $|1\rangle$ in the HHL ancilla results in a quantum state $|\tilde{x}\rangle$ in the state register, which is proportional to the required $\vec{x}$. For any application, where some feature of $\vec{x}$ is to be evaluated, an additional module can be appended to the HHL circuit. For example, if the application of interest to us is quantum chemistry and we intend to compute molecular energies using HHL, then a controlled-swap test circuit module could be added at the end of the HHL circuit, as discussed in Reference [adaptHHL].
Figure 1: Schematic representation of the HHL algorithm.
@@ III HHL with qutrits
In this section, we discuss the HHL algorithm using qutrits in detail.
@@ III.1 A brief introduction to ternary logic
We begin our discussion with a concise introduction to ternary logic. Ternary logic is a multi-valued logic system that extends the binary system of logic by adding a third state ‘$2$’ to $\{0,1\}$. Thus, the computational basis for a ternary quantum state would be $\{|0\rangle,|1\rangle,|2\rangle\}$. A general qutrit state can be defined as $|\psi\rangle=\alpha_{0}|0\rangle+\alpha_{1}|1\rangle+\alpha_{2}|2\rangle$ where $|\alpha_{0}|^{2}+|\alpha_{1}|^{2}+|\alpha_{2}|^{2}=1$. Similarly, an $n$-qutrit state would have $3^{n}$ different possible states. In terms of the state vector notation, we can represent the computational basis vectors as -
@@ III.2 Basic gates for qutrits
In this study, we work with a palette of one- and two-qutrit gates. This includes the one- and two-qutrit Clifford gates introduced in Reference [gottesman_gates], that is, the qutrit equivalents of the Pauli X and Z gates, and the CX gate. We also pick the H gate, the phase gate, and planar rotation gate, along with their controlled counterparts of the latter two [qudit-algorithms]. Our choice of gates to build the qutrit HHL circuit is motivated by convenience in circuit construction. Thus, we describe in detail only those gates.
III.2.1 One-qutrit gates
@@
    Increment gate (X): The qutrit X gate increments a computational basis state of the qutrit in cyclic order. It is specified in its matrix representation as
    $X=\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{pmatrix}$.
    In the outer product notation,
    $X=|1\rangle\langle 0|+|2\rangle\langle 1|+|0\rangle\langle 2|.$
    The action of this gate can be described as follows -
    $X|0\rangle$ $=|1\rangle$
    $X|1\rangle$ $=|2\rangle$
    $X|2\rangle$ $=|0\rangle.$
  2. 2.
    Z gate:
    $Z=\begin{pmatrix}1&0&0\\ 0&\omega&0\\ 0&0&\omega^{2}\end{pmatrix}$.
    In the outer product notation,
    $Z=|0\rangle\langle 0|+\omega|1\rangle\langle 1|+\omega^{2}|2\rangle\langle 2|,$
    where $\omega=e^{\frac{2\pi i}{3}}$ is the third root of unity.
    The action of this gate can be described as follows -
    $Z|0\rangle$ $=|0\rangle$
    $Z|1\rangle$ $=\omega|1\rangle$
    $Z|2\rangle$ $=\omega^{2}|2\rangle.$
  3. 3.
@@
    $H=\frac{1}{\sqrt{3}}\begin{pmatrix}1&1&1\\ 1&\omega&\omega^{2}\\ 1&\omega^{2}&\omega\end{pmatrix}$.
    In the outer product notation,
    $H=\frac{1}{\sqrt{3}}\sum_{j,k\;\epsilon\;\{0,1,2\}}\omega^{j\cdot k}|j\rangle\langle k|.$
@@
  4. 4.
@@ III.2.2 Two-qutrit gates
    Controlled increment (CX): The controlled increment gate increments the state of the target qutrit depending on the control qutrit. This gate is also called the SUM gate as the operation is equivalent to a modulo 3 addition. The $CX$ gate is defined as -
@@
    where $|k\rangle$ is any qutrit state, and the action of the $X$ gate is already defined in the list of single qutrit gates.
@@
    Controlled $P_{l}$ gate: The action of the controlled $P_{l}$ gate is described as -
@@
    Generalized controlled $R_{ij}(\theta)$: In this construction, $R_{ij}(\theta)$ gate acts on the target qutrit only if the state of the control qutrit is equal to $|m\rangle$, where $m$ is either 0, 1, or 2.
@@ III.3 Qutrit HHL: working principles and quantum circuit
The HHL algorithm has been described briefly in Section II. In this section, we discuss our extension to the qutrit framework. The general problem statement remains the same, that is, solving a system of linear equations of the form $A\vec{x}=\vec{b}$. The $\vec{b}$ of dimension $N\crossproduct 1$ is encoded in a state register containing $m$ qutrits using amplitude encoding such that $3^{m}=N$. In addition to the state register, we have a clock register comprising of $n_{r}$ qutrits. These are useful for capturing the eigenvalues via the QPE module. Lastly, we have a single qutrit ancilla, also known as the HHL ancilla, which will be used for eigenvalue inversion through the controlled rotation module.
@@
We begin by discussing the quantum Fourier transform (QFT) module, since its inverse (IQFT) occurs in the QPE module. Figure 2 gives the circuit for QFT. Although the circuit is similar to that for qubits, the gate definitions are modified in the qutrit framework. We define the generalized $P_{l}$ gate for qutrits as specified in Section III.2. Implementing the QFT circuit on a general qutrit state $|j_{1},j_{2},\ldots j_{n}\rangle$, we obtain -
@@
The circuit for the QPE module is given in Figure 3. QPE consists of mainly controlled unitaries, with the unitary matrix of the form $U=e^{iAt}$ where $A$ is the Hermitian matrix that occurs the system of linear equations to be solved, and $t$ is a problem-dependent constant such that $U(\sum_{i=0}^{3^{m}-1}b_{i}|u_{i}\rangle)=\sum_{i=0}^{3^{m}-1}b_{i}e^{2\pi i\phi_{i}}|u_{i}\rangle$ where $\phi_{i}$s $\in[0,1]$ are the eigenvalues of $A$. The clock registers are initially all in state $|0\rangle$. First, all the clock registers are acted upon by the ternary Hadamard gates resulting in -
@@
Upon the application of the series of controlled unitaries shown in Figure 3, we obtain
@@
Substituting $\phi_{i}$s in the form $\tilde{\phi_{i}}=0.j_{1}j_{2}\ldots j_{n_{r}}^{(i)}$, we obtain an expression equivalent to that of QFT, and hence the inverse QFT module takes the eigenvalues of $A$ from the phase to the qutrit state. The precision is, of course, limited by the number of clock register qutrits used and the resultant state after the QPE module $|\psi_{C}\rangle=\sum_{i=0}^{3^{m}-1}b_{i}\tilde{|\phi_{i}\rangle}\otimes|u_{i}\rangle_{m}$. Reference [qudit-algorithms] briefly describes the QFT and QPE modules for higher-dimensional qudits, in general. We have worked them out in detail for the specific case of qutrits.
Figure 4: Uniform controlled-rotation circuit for 1 control qutrit.
Next, we have the parametrized controlled-rotation module, which is used for eigenvalue inversion. Rotations in general could be quite complex in a qutrit framework, and for the purposes of our HHL algorithm, we choose to use planar rotations, that is, a rotation of any two of the three orthogonal qutrit basis states. Without loss of generality, we consider $|0\rangle$ and $|1\rangle$ for our rotations (any other combination can be used equivalently). The rotation gate as well as the controlled-rotation gates have been described in detail in Section III.2. The simple example circuit shown in Figure 4 serves as an illustration of the workings of uniformly controlled rotations. We assume that the control qutrit is in a general state $|\psi\rangle=\alpha_{0}|0\rangle+\alpha_{1}|1\rangle+\alpha_{2}|2\rangle$. The action of the three controlled-rotation gates (UCR) can be described in the outer product notation as $|0\rangle\langle 0|\otimes R_{01}(\theta_{0})+|1\rangle\langle 1|\otimes R_{01}(\theta_{1})+|2\rangle\langle 2|\otimes R_{01}(\theta_{2})$. Therefore,

### v2

@@ Abstract
We extend the Harrow-Hassidim-Lloyd (HHL) algorithm, which is well-studied in the qubit framework, to its qutrit counterpart (which we call qutrit HHL, as opposed to qubit HHL, which is HHL using qubits), and develop a program for its implementation. We design Weyl-Heisenberg gadgets, the qutrit equivalents of Pauli gadgets, and come up with a practical implementation scheme for qutrit HHL. We test HHL with qutrits for simple matrices and verify the results against the expected outcomes. We apply the algorithm to quantum chemistry, and in particular, to the potential energy curve calculations of the model problem of the Hydrogen molecule in the split valence basis. We do so for two cases: 1-qutrit and 2-qutrit input states, where the latter makes use of our gadgets. We compare the number of qudits and the number of gates required between qubit and qutrit HHL implementations. In general, we find that for a fixed precision, the qutrit HHL circuit requires fewer number of qudits and comparable number of two-qudit gates than its qubit counterpart.
@@ I Introduction
All of the above discussed works involve qubits. Recent works have identified that using qudits for quantum algorithms can have potential advantages over their qubit counterparts [qudit-adv, qutrit-benchmarking]. Several algorithms like Deutsch-Josza, Bernstein-Vazirani, quantum Fourier transform (QFT), QPE, etc, have been discussed in the qudit framework [qudit-algorithms]. Despite the large volume of ongoing research in quantum algorithms and computing, there remains a relative paucity of literature on higher-dimensional qudits, and in particular, qutrits, where a qutrit has an additional quantum level relative to a qubit, but which is a natural extension to consider. Furthermore, in view of some of the recent advances in qutrit hardware [qutrit-hardware, qutrit-hardware-gates, qutrit-benchmarking], it is timely to explore the domain of quantum algorithms with qutrits. Therefore, given the potential of HHL for offering advantage across a wide variety of problems and the timely need to investigate quantum algorithms using qutrits, we extend the HHL algorithm to the qutrit framework (which we shall term as qutrit HHL hereafter).
We discuss a scheme for the practical implementation of qutrit HHL, given a simple gate set comprising of 1- and 2-qutrit gates. To this end, we come up with qutrit equivalents of Pauli gadgets [pauli_gadgets], which we call Weyl-Heisenberg (WH) gadgets, to implement unitaries expressed as matrix exponentials. We also derive relations for qutrit operators, which include expressing any $3^{n}\times 3^{n}$ matrix (an $n$- qutrit matrix) as a linear combination of tensor products of the nine 11- qutrit WH operators and an efficient circuit construction for the controlled-WH gadgets.
Furthermore, given the potential of applying HHL to quantum chemistry, we consider it as a pilot application for qutrit HHL, and to that end, carry out calculations to generate the potential energy curve for the model system of $H_{2}$ in a split valence basis. This is followed by a comparative analysis of the costs involved between qubit and qutrit HHL calculations. To the best of our knowledge, one previous work proposes qutrit HHL, but only presents a preliminary circuit design [old-qutrit-hhl].
The subsequent sections are organized as follows: Section II introduces the HHL algorithm for qubits. Following a primer on ternary logic and qutrit gates in Sections III.1 and III.2 respectively, we move to our qutrit HHL implementation, which is discussed in Section III.3. Next, in Section IV, we discuss in detail the decomposition of multi-qutrit unitaries to 1- and 2-qutrit gates, along with the step by step scheme to practically implement the qutrit HHL as a quantum circuit. We then provide a brief introduction to the quantum chemistry application, in which we solve for the correlation energy of a molecule using HHL (Section V). We then discuss results from our numerical simulations in Section VI, which include toy matrices (Section VI.1) and a quantum chemistry application (Section VI.2). We then carry out a comparative analysis between qubits and qutrits in Section VII, where we discuss estimates for the number of qudits required (Section VII.1) and the number of 2-qudit gates incurred for HHL calculations in the qubit and qutrit cases (Section VII.2). Lastly, in Section VII.3 we discuss the gate counts for implementation of physical Hamiltonians, comparing both the HHL implementations. We finally conclude in Section VIII.
@@ II The HHL algorithm
The HHL algorithm is discussed in detail, for example, in Reference [HHL_walkthru], but here we provide only the relevant details that we would need for a subsequent qutrit extension of the algorithm. Figure 1 gives a schematic quantum circuit representation of the algorithm. The $N\crossproduct 1$ dimensional vector, $\vec{b}$, which is an input to the algorithm, is encoded in an $m$-dimensional normalized quantum state $|b\rangle_{m}$ in the state register via amplitude encoding such that $2^{m}=N$ for qubits. On the other hand, the matrix $A$ is used in the circuit as a part of a controlled-unitary of the form $C\ U=C\ e^{iAt}$, where $U=e^{iAt}$ for an appropriately chosen scalar, $t$. As the figure shows, the HHL circuit broadly comprises of three parts - (a) QPE, (b) Controlled-rotation, and (c) Inverse of QPE. These steps are followed by measurement and post-selection. The above-mentioned controlled-unitaries, which are controlled on a set of $n_{r}$ clock register qubits, are used for estimating the phase. At the end of the QPE module, the eigenvalues of $A$ are stored in the clock register with a precision of $n_{r}$ digits. The parametrized controlled-rotation module, $CR(\theta)$, implements inversion of eigenvalues using an ancillary qubit that we call the HHL ancilla for brevity. The inverse QPE module is then applied in order to disentangle the clock register from the rest of the system. Post-selection of $|1\rangle$ in the HHL ancilla results in a quantum state $|\tilde{x}\rangle$ in the state register, which is proportional to the required $\vec{x}$. For any application, where some feature of $\vec{x}$ is to be evaluated, an additional module can be appended to the HHL circuit. For example, if the application of interest to us is quantum chemistry and we intend to compute molecular energies using HHL, then a controlled-swap test circuit module could be added at the end of the HHL circuit, as discussed in Reference [adaptHHL].
Figure 1: Schematic quantum circuit representation of the HHL algorithm.
@@ III HHL with qutrits
In this section, we discuss in detail the HHL algorithm using qutrits.
@@ III.1 A brief introduction to ternary logic
We begin our discussion with a concise introduction to ternary logic. Ternary logic is a multi-valued logic system that extends the binary system of logic by adding a third state ‘$2$’ to $\{0,1\}$. Thus, the computational basis for a ternary quantum state would be $\{|0\rangle,|1\rangle,|2\rangle\}$. A general 1-qutrit state can be defined as $|\psi\rangle=\alpha_{0}|0\rangle+\alpha_{1}|1\rangle+\alpha_{2}|2\rangle$ where $|\alpha_{0}|^{2}+|\alpha_{1}|^{2}+|\alpha_{2}|^{2}=1$. Similarly, an $n$-qutrit state would have $3^{n}$ different possible basis states. In terms of the state vector notation, we can represent the 1-qutrit computational basis vectors as -
@@ III.2 Basic gates for qutrits
In the ternary framework, logical operators are extended to $3\crossproduct 3$ unitaries. In this case, the regular definitions of the qubit gates are not retained, and need to be modified. Just as the Pauli matrices $\sigma^{x}$, $\sigma^{y}$ and $\sigma^{z}$, along with identity form a complete basis for the space of operators acting on the Hilbert space for a qubit, we have the WH operators (Table 1), which include identity, forming the basis for operators acting on qutrits [WH_op_ref]. They are defined as -
$X^{a}Z^{b}\;\;\;\;\;\;\;\;\forall a,b\in\{0,1,2\}.$
Table 1: Weyl–Heisenberg operators for a qutrit, where $\omega=e^{2\pi i/3}$.
$(a,b)$ $(0,0)$ $(1,0)$ $(2,0)$ $(0,1)$ $(0,2)$ $(1,1)$ $(1,2)$ $(2,1)$ $(2,2)$
$W_{a,b}$ II $X$ $X^{2}$ $Z$ $Z^{2}$ XZXZ $XZ^{2}$ $X^{2}Z$ $X^{2}Z^{2}$
Matrix form $\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}$ $\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{pmatrix}$ $\begin{pmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{pmatrix}$ $\begin{pmatrix}1&0&0\\ 0&\omega&0\\ 0&0&\omega^{2}\end{pmatrix}$ $\begin{pmatrix}1&0&0\\ 0&\omega^{2}&0\\ 0&0&\omega\end{pmatrix}$ $\begin{pmatrix}0&0&\omega^{2}\\ 1&0&0\\ 0&\omega&0\end{pmatrix}$ $\begin{pmatrix}0&0&\omega\\ 1&0&0\\ 0&\omega^{2}&0\end{pmatrix}$ $\begin{pmatrix}0&\omega&0\\ 0&0&\omega^{2}\\ 1&0&0\end{pmatrix}$ $\begin{pmatrix}0&\omega^{2}&0\\ 0&0&\omega\\ 1&0&0\end{pmatrix}$
Some other gates that are used in this work are as follows:
III.2.1 1-qutrit gates
@@
    $H=\frac{1}{\sqrt{3}}\begin{pmatrix}1&1&1\\ 1&\omega&\omega^{2}\\ 1&\omega^{2}&\omega\end{pmatrix}$,
    where $\omega=e^{2\pi i/3}$. In the outer product notation,
    $H=\frac{1}{\sqrt{3}}\sum_{j,k\;\in\;\{0,1,2\}}\omega^{j\cdot k}|j\rangle\langle k|.$
@@
  2. 2.
    S gate:
    $S=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&\omega\end{pmatrix}$.
    The action of this gate can be described as follows -
    $S|0\rangle$ $=|0\rangle$
    $S|1\rangle$ $=|1\rangle$
    $S|2\rangle$ $=\omega|2\rangle.$
  3. 3.
@@
    Controlled-increment ($CX$): The controlled-increment gate increments the state of the target qutrit depending on the control qutrit. This gate is also called the SUM gate as the operation is equivalent to a modulo 3 addition. The $CX$ gate is defined as -
@@
    where $|k\rangle$ is any 1-qutrit state, and the action of the $X$ gate is already defined in the list of 1-qutrit gates.
@@
    Controlled-$P_{l}$ gate: The action of the controlled-$P_{l}$ gate is described as -
@@
    Generalized controlled-$R_{ij}(\theta)$: In this construction, $R_{ij}(\theta)$ gate acts on the target qutrit only if the state of the control qutrit is equal to $|m\rangle$, where $m$ is either 0, 1, or 2.
@@ III.3 Qutrit HHL: working principles and quantum circuit
The HHL algorithm has been described briefly in Section II. In this section, we discuss our extension to the qutrit framework. The general problem statement remains the same, that is, solving a system of linear equations of the form $A\vec{x}=\vec{b}$. The $\vec{b}$ of dimension $N\crossproduct 1$ is encoded in a state register containing $m$ qutrits using amplitude encoding such that $3^{m}=N$. In addition to the state register, we have a clock register comprising of $n_{r}$ qutrits. These are useful for capturing the eigenvalues via the QPE module. Lastly, we have a single qutrit ancilla, also known as the HHL ancilla, which will be used for eigenvalue inversion through the controlled-rotation module.
@@
We begin by discussing the QFT module, since its inverse (IQFT) occurs in the QPE module. Figure 2 gives the circuit for QFT. Although the circuit is similar to that for qubits, the gate definitions are modified in the qutrit framework. We define the generalized $P_{l}$ gate for qutrits as specified in Section III.2. Implementing the QFT circuit on a general n-qutrit state $|j_{1},j_{2},\ldots j_{n}\rangle$, we obtain -
@@
The circuit for the QPE module is given in Figure 3. QPE consists of mainly controlled-unitaries, with the unitary matrix of the form $U=e^{iAt}$ where $A$ is the Hermitian matrix that occurs the system of linear equations to be solved, and $t$ is a problem-dependent constant such that $U(\sum_{i=0}^{3^{m}-1}b_{i}|u_{i}\rangle)=\sum_{i=0}^{3^{m}-1}b_{i}e^{2\pi i\phi_{i}}|u_{i}\rangle$ where $\phi_{i}$ $\in[0,1]$ are the eigenvalues of $A$. The clock registers are initially all in state $|0\rangle$. First, all the clock registers are acted upon by the ternary Hadamard gates resulting in -
@@
Upon the application of the series of controlled-unitaries shown in Figure 3, we obtain
@@
Substituting $\phi_{i}$ in the form $\tilde{\phi_{i}}=0.j_{1}j_{2}\ldots j_{n_{r}}^{(i)}$, we obtain an expression equivalent to that of QFT, and hence the inverse QFT module takes the eigenvalues of $A$ from the phase to the qutrit state. The precision is, of course, limited by the number of clock register qutrits used and the resultant state after the QPE module $|\psi_{C}\rangle=\sum_{i=0}^{3^{m}-1}b_{i}\tilde{|\phi_{i}\rangle}\otimes|u_{i}\rangle_{m}$. Reference [qudit-algorithms] briefly describes the QFT and QPE modules for higher-dimensional qudits, in general. We have worked them out in detail for the specific case of qutrits.
Figure 4: Uniformly controlled-rotation circuit for 1 control qutrit.
Next, we have the parametrized controlled-rotation module, which is used for eigenvalue inversion. Rotations in general could be quite complex in a qutrit framework, and for the purposes of our HHL algorithm, we choose to use planar rotations, that is, a rotation of any two of the three orthogonal qutrit basis states. Without loss of generality, we consider $|0\rangle$ and $|1\rangle$ for our rotations (any other combination can be used equivalently). The rotation gate as well as the controlled-rotation gates have been described in detail in Section III.2. The simple example circuit shown in Figure 4 serves as an illustration of the workings of uniformly controlled-rotations (UCR). We assume that the control qutrit is in a general state $|\psi\rangle=\alpha_{0}|0\rangle+\alpha_{1}|1\rangle+\alpha_{2}|2\rangle$. The action of the three UCR gates can be described in the outer product notation as $|0\rangle\langle 0|\otimes R_{01}(\theta_{0})+|1\rangle\langle 1|\otimes R_{01}(\theta_{1})+|2\rangle\langle 2|\otimes R_{01}(\theta_{2})$. Therefore,